Opened 9 years ago

Closed 3 years ago

#333 closed flight processing (fixed)

CEH07/04, flight day 102b/2010, Moorhouse

Reported by: benj Owned by: knpa
Priority: alpha 4 low Milestone: 2010 data processing completion
Component: Archiving Keywords:
Cc: Other processors:

Description (last modified by emca)

Data location: ~arsf/arsf_data/2010/flight_data/uk/CEH07_04-2010_102b_Moorhouse

Data arrived from ARSF via network transfer on 12th April 2010.

Scientific objective: Unknown (can't find application)

Priority: a4l, process hyperspectral as a3

PI: Andrew Wilson
Also Clare Rowland at CEH Lancaster (see contact details page for email)

Collected as a single collection with day 102a (Hellifield, #332) - information on the same logsheet, combined nav data, etc. Have split into two separate collections for processing.

Sensors:

  • Eagle
  • Hawk
  • Leica LIDAR (08/10/2010)
  • Camera (01/09/2010)

Change History (26)

comment:1 Changed 9 years ago by benj

  • Description modified (diff)

Gary says priority is a4 from before they were recorded as high/med/low. Process as a4l for now, since original PI has retired, we've had no other contacts chasing this and that way it won't push our processing schedule back.

comment:2 Changed 9 years ago by benj

  • Description modified (diff)

Andrew has asked that we send him the raw hyperspectral data for this flight and he'll process them. We just need to do the applanix and the Lidar.

Andrew should send us back a copy of the data for archiving when he's done, but in case he doesn't I'm leaving it in the processing queue, it'll just stay at the bottom. If/when Andrew does return the processed data to us it can come off the queue.

comment:3 Changed 9 years ago by pmlvis12

IPAS navigation processing done.

comment:4 Changed 9 years ago by ella

POSProc failed to make an Sbet file for this project during applanix processing, when attempt was made error message read too many consecutive IIN measurment rejections. Needs someone to take a look.

comment:5 Changed 9 years ago by jb362

  • Description modified (diff)

Started camera processing

comment:6 Changed 9 years ago by jb362

Camera awaiting delivery check.

comment:7 Changed 9 years ago by ella

Camera delivery check complete. Everything looks good; ready to deliver.

comment:8 Changed 9 years ago by jb362

Camera data delivered 01/09/2010.

comment:9 Changed 9 years ago by jb362

  • Description modified (diff)

comment:10 Changed 9 years ago by emca

Starting LiDAR processing

comment:11 Changed 8 years ago by emca

LiDAR lines Moor2 and Moor4 had incorrect range settings during data acquisition resulting a hole in the data for these lines.

LiDAR data is ready for delivery check at
~airborne/workspace/CEH07_04-2010_102b_Moorhouse/delivery/20100930/CEH07_04

comment:12 Changed 8 years ago by benj

  • Description modified (diff)

comment:13 Changed 8 years ago by iopa

Delivery checking LiDAR.

comment:14 Changed 8 years ago by iopa

Delivery check complete, LiDAR ready to deliver.

comment:15 Changed 8 years ago by emca

  • Description modified (diff)

LiDAR data dispatched to A Wilson and C Rowland 08/10/2010

comment:16 Changed 8 years ago by emca

Camera data sent to C Rowland 15/10/2010

comment:17 Changed 8 years ago by emca

Starting hyperspectal processing

comment:18 Changed 7 years ago by emca

  • Component changed from Processing: general to Archiving
  • Owner set to benj

Hyperspectral processing with az was not finished and we didn't receive a copy of the data processed by Andrew Wilson.

Claire Rowland @ CEH emailed 11/05/2012 asking about the Eagle and Hawk data.

The data was reprocessed using APL by adbe and made available to Claire via ftp 30/05/2012

Ready for archiving

comment:19 Changed 7 years ago by knpa

NEODC have had a request for this project from Mark Kincey @ Durham University.

Will combine reprocessed hyperspectral with rest of project and archive so it can be accessed.

comment:20 Changed 7 years ago by knpa

  • Owner changed from benj to knpa
  • Status changed from new to assigned

comment:21 Changed 7 years ago by knpa

Finished preparing for archiving, made available to NEODC.

comment:22 Changed 7 years ago by knpa

Note that this project has been combined with the reprocessed hyperspectral.

comment:23 Changed 7 years ago by knpa

NEODC have archived it.

Moved to archive disk.

comment:24 Changed 6 years ago by knpa

Removed workspace version (original processing).

Leaving ticket open until reprocessing is archived.

comment:25 Changed 3 years ago by dac

Archiving

Archived data available from NEODC (http://browse.ceda.ac.uk/browse/neodc/arsf/2010/CEH07_04/CEH07_04-2010_102b_Moorhouse/) started uploading raw data.

comment:26 Changed 3 years ago by dac

  • Resolution set to fixed
  • Status changed from assigned to closed

Archiving

Raw data also available from NEODC. Marking as closed.

Note: See TracTickets for help on using tickets.